OpenStudy (anonymous):

How do I find the indefinite integral of: ∫(√x) + (1/(4√x))

OpenStudy (anonymous):

I got 2/3(x^3/2)+2(x^1/2)+C however it is stated to be wrong

OpenStudy (kinggeorge):

Can you show me how you got to that result?

OpenStudy (kinggeorge):

P.S. It's almost correct.

OpenStudy (anonymous):

Sure x^1/2 + 1/4 x^-1/2 (x^3/2)/(3/2)+(1/4)(x^1/2)/(1/2)

OpenStudy (anonymous):

Am I missing a step?

OpenStudy (kinggeorge):

That looks good to me. Now what is \[\Large \frac{1/4}{1/2}?\]

OpenStudy (anonymous):

2

OpenStudy (kinggeorge):

Almost. \[\Large \frac{1/4}{1/2}=\frac{2}{4}=\frac{1}{2}\]In general, \[\Large \frac{a/b}{c/d}=\frac{ad}{bc}\]

OpenStudy (anonymous):

ah careless mistake

OpenStudy (anonymous):

so then instead of the 2√x its 1/2√x?

OpenStudy (kinggeorge):

That looks right to me.

OpenStudy (anonymous):

ok 2/3(x^3/2)+1/2(x^1/2) is the correct answer?

OpenStudy (kinggeorge):

With a +C at the end.

OpenStudy (anonymous):

Always forget the C, thanks a lot you were a really big help. You truly are the king.

OpenStudy (kinggeorge):

You're welcome, and thanks :P